# Application of AI to formal methods — an analysis of current trends

### v1

@@ 5.2 RQ 2: Content type of contributions
Lastly, we investigated the found data sets. Here, multiple observations took place.
First, we searched for high-quality data sets, enabling other researchers to conduct their research efficiently. As pointed out earlier, we found 21 data sets. In Fig. 14 (see also Appendix B), we see the respective scopes of the data sets. Of the 21 data sets, 15 focused on TP, while 3 aimed at synthesis, 1 at SAT, 1 at SMT, and 1 at MC.
However, we were also interested in the data sets, which were not novel contributions but instead found reuse in our found studies. For this, we skimmed all our studies for data set mentions. As it turns out, of the 189 articles, 124 used external data sets. The 65 remaining studies generate random samples or do not mention their data source. Noteworthy is that when random data is used, the generated data is seldomly shared, so the reproduction of results is hindered.
The datasets found for the 124 contributions are largely heterogenic. This means that most studies utilized different datasets. However, we could find some data repositories that had multiple usages. The Mizar Library [^6] https://mizar.uwb.edu.pl/library/ (and associated libraries like, e.g., MPTP[47]) saw 23 usages, TPTP Problem Library [^7] https://tptp.org/TPTP/ has 7 mentions, The ProB machine library [^8] https://prob.hhu.de/w/index.php?title=Download got 5 mentions, HoL [^9] https://isabelle.in.tum.de/library/ 5, SMTLib [^10] https://smt-lib.org 5, SATComp [^11] https://satcompetition.github.io 5, SATLib [^12] https://www.cs.ubc.ca/~hoos/SATLIB/benchm.html 4, CoqGym [^13] https://github.com/princeton-vl/CoqGym 4. The full list is available in the repository linked in Section 1.
@@ 6 Discussion
In the following, we will evaluate the results and conclude to answer our research questions. Further, we will outline lessons learned to provide actionable insights for conducting future studies in the area.
@@ 6.1 RQ 1: Demographics of the research area
For RQ 1.1, we see a general growth in the research field. However, from the presented data points it cannot be inferred alone whether this trend follows the overall trend of AI’s rise in popularity or if it is independent of it.
For RQ 1.2, we see a strong focus on conference papers, short papers, and workshops. Together with the singular book chapter, this indicates an emerging field, as there is a tendency to rely on publication methods supporting short review cycles. This is further supported by the fact that the only book chapter was a case study. Contrasting this with RQ 1.1, we assume that the research interest will at least not plummet in the foreseeable future but might plateau for the time being.
For RQ 1.3, we can see that the bulk of works lay within the area of methodologies. The reason for this could be the current paper requirements, requesting a novel approach and some evaluation within the paper. Another reason might be AI’s applicative nature, which often allows it to serve as a support system to automate parts of the work. We also want to highlight the possibility that we primarily introduced methodology papers based on the nature of our search query, which focuses on the application of AI in the FM domain. However, due to rigorous snowballing, we are confident we correctly assessed the state of research in this field.
What is more interesting is what we do not see: many case studies and benchmarks. Recalling Fig. 6, we only see two publications providing case studies and two providing benchmarks. This may hint at a further gap in maturity, especially as these contribution types are primarily present in journal and book contributions. One would assume more such publications if a set of robust solutions for common problems were readily available so that researchers could compare their effectiveness against each other. It also highlights the need for proper and uniform benchmark sets within the community which can be shared between different research groups.
@@ 6.2 RQ 2: Content type of contributions
In RQ 2, we were interested in how AI application is distributed in FM. For this, we posed RQs 2.1, 2.4, 2.2 and 2.3 to investigate which AI techniques find appreciation in the community, which FM areas already are subject to AI application, if certain AI techniques are predominantly used in specific FM fields, and whether there exist publicly available data sets which can be used in further research.
@@
For RQ 2.1, we saw in Fig. 7 Most studies use NNs, followed by RL, and NLP and EAs are far behind. From a paradigmatic point of view, it makes sense to see these four on top, as they all attempt to solve a different problem setting. NNs are a stand-in for classification and regression tasks, RL is applied for learning behavioral policies, NLP is processing texts resembling natural language or code by extension, and EAs excel in optimization problems. However, it is interesting to see NNs being so far ahead of the rest with 83 publications utilizing them. Especially, since they are not the only algorithms suitable for classification and regression tasks, but competing algorithms such as KNN, SVM, or decision trees are, at least as solely applied AI techniques, vastly underrepresented. These algorithms, which we categorized as classical approaches in Section 2.3, are, however, not forgotten, as we can see in publications which contrast multiple different AI algorithms (Fig. 7(b)). This focus on NNs when only one technique is used can have multiple underlying reasons. First, implementation, training, and fine-tuning of NNs is a time-consuming task, and might take up the resources needed to investigate further algorithms. Second, we have seen in Fig. 3 that the research area gained traction in the early 2010s. As this coincides with the advent of deep learning [48], this might suggest the need for NNs to overcome the complexities the area of FM offers, and that more classical ML algorithms are not suitable for most tasks. Third, there might exist a bias towards using NNs from the get-go in the community instead of applying more straightforward techniques first.
@@
However, the prominence of more classical algorithms in the case of multiple AI techniques showcases that the community seems to be aware that other algorithms might be suitable for their problems as well. This is especially amplified by Fig. 8 showing that most frequently, it is methodologies that rely on various techniques to solve problems.
@@
Regarding maturity, Fig. 9 indicates no noteworthy increase in techniques besides NLP. For 2023, we established insufficient data points, so we may see a rise in NLP-based contributions, which would fit a more general advent of NLP-related contributions due to the rise of transformer architectures and LLMs.
@@
For RQ 2.2, we have seen in Fig. 10(a) that TP and SAT are the predominant FM areas in which AI is applied, making up 68.8 % of all publications in our five-year corpus. We explain this by the fact that both areas are also of interest to non-FM communities. A detailed view of the authors’ origin domains would be necessary to verify this assumption. However, this is still beneficial for the FM community as a whole. For instance, Fig. 11(a) shows a strong focus on FM-relevant daily use cases. Figure 11(b) reveals much current work regarding first-order logic on the especially relevant FM topics of premise selection. While this is a good trend, complex FM models rely on complicated data types and require sophisticated reasoning.
@@
In regards to SAT, we can take away from Fig. 10(c) that much focus lies on SAT solving directly, while other authors also attempt to predict satisfiability. The latter seems to stand in contradiction with the rigorousness needed in FM, as the probabilistic nature of AI approaches lacks the certainty of the predicted satisfiability as a whole. This can be a product of cross-fertilization, as SAT is also relevant for non-FM communities.
@@
Figure 10(d) shows that the interest in synthesis is low compared to other topics and primarily focuses on model generation. Given the recent rise of generative AI approaches, we see room for promising research in this area in the future.
@@
A similar low interest seems to be in model checking. Figure 12(b) shows a relatively low number of contributions in that area. The direct comparison with the TP field is interesting, as MC also contributes very little in tool and tool contributions, hinting at less application-focused research. Figure 12(a) shows a relatively consistent interest in recent years, while Fig. 12(b) emphasizes how little is done for current practitioners in terms of tool (improvements), benchmarks, training data, and case studies.
@@
For RQ 2.3, we saw in Fig. 13(a) that NNS is the undeniable favorite in almost all FM areas investigated. We already discussed why NNs might be more popular than other algorithms suitable for the same problem tasks above for RQ 2.1. However, this suggests again how the FM community mainly focuses on classification or regression problems to solve rather than execution policies (RL) or direct processing of texts or code (NLP). The slight exception we see is in the area of model checking, which focuses most on EAs, suggesting a more substantial prevalence of optimization problems in this area. From building the larger corpus of 457 studies, we know, however, from sighting various titles that we missed many EA papers due to our 5-year restriction. Presumably, EA was more prominently used before deep learning took off.
@@
Finally, for RQ 2.4, we found 19 data set (or training environment) contributions in our corpus of 189 studies. Considering the two separate, data set-only publications which did not pass IC 1 (see Section 4.5), this leaves us with 21 total data set contributions.
Together with the analyzed high heterogeneity of the used datasets, as already mentioned in Section 5.1.3, there may be a general underlying issue. There seems to be no defined benchmark against which to run new insights. While datasets are publicly available for TP and SAT solving, the number of usable, established, and available benchmark data sets is minuscule for other areas.
@@
Concluding, we see a strong need for unified defaults to not only allow reproducibility but also to introduce familiar benchmarks into the field. Possible issues hindering such a development might be found in the various existing formalisms, which are not necessarily used or even supported by individual research groups. Such boundaries, of course, weaken the impact a data set publication might have onto researchers working with differing formalisms.
@@ 6.3 Suggested research directions
As pointed out in our discussion to RQ 2.4, we are concerned with the lack of established data sets and benchmarks in the FM su-communities. While for TP, a good amount of contributions exist already tackling this problem, the other areas seemingly fall short. This creates room for reproducibility issues between research groups and hinders a more precise comparison between results as established benchmarks could provide.
Developing strong baselines might not only increase comparability between different methods but might also reduce the entry hurdle for new researchers due to publicly available data sets. The main challenge is to create data sets which can be used for different formalisms. This would allow smaller communities to still benefit from available data which might be posed in a more popular formalism while furthering research on their area.
An orientation on how such a unified benchmark environment could look like would be OpenAIGym [50], a Toolkit for RL to foster research by enhancing reproducibility.
@@
In the age of big data, we were expecting more studies that apply data mining techniques, be it for feature engineering or performance analytical approaches. This was not the case. While it might very well be that the kind of problems solvable by data mining are not relevant to the FM communities, it is also possible that potential benefits are not well-known.
We propose to research if and how data mining approaches can benefit FM. First, this would highlight potential benefits to the community. Second, due to the investigative nature of data mining, this might strengthen our insides in our tools and best practices. Consequently, this might further the quality of other AI applications in FM due to new knowledge.
@@
While LLMs, at this point, are established tools that found their way into daily applications, we only found little use of them so far in our corpus of studies. This can be due to their relatively high employment costs. Another reason could be that respective studies are still in the process of being published and were hence invisible to our systematic mapping study. Nevertheless, we see strong potential for generative AI approaches, especially in the areas of synthesis and auto-formalization. However, also simple test-case generation or the generation of documentation for existing models seem promising starting points to build trust in the process.
We envision substantial benefits from applying generative AI during FM development and point out the need for more research. One can easily envision tool chains of automatic generation, checking, and documentation of formal models with reproducible and provable results. However, we also see the need for more exploration regarding how generative AI can be helpful besides text production. For instance, how could generative AI be leveraged in proof mining? What stages in the formal development process can benefit from generated inputs, and what types could and should be generated? Pure generation of code and text might not be the ceiling for FM.
Increase the potentials of AI in model checking
Another underrepresented topic is model checking. Anecdotally, we know model checking had much EA applied in the past while it was a seemingly unpopular application area for AI in our five-year corpus. However, we see a significant potential to revisit respective research with a modern lens. Especially in the context of bounded model checking (BMC) or statistical model checking (SMC) and faulty state prediction. BMC cuts off parts of a model’s state space for performance and computability. SMC checks states until a certain confidence in the absence of faulty states is reached. Both approaches already do not fully traverse the state space and can only indicate the absence of faults.
AI-enforced model checking could alleviate BMC, SMC, and model checking as a whole by predicting potential faulty states or needed search depths, learning search policies via RL to reach faulty states in a more targeted manner, or predict equivalence classes of states for search space reduction.
@@ 7 Threats to validity
A study of this type and extent can only be conducted with the possibility of encountering threats that threaten validity. We orient ourselves at Zhou et al. Zhou et al. [30] for addressing the most critical ones.
@@
A potential threat to the conclusion cloud is our selection of articles and our data source. To ensure the quality of our data, we carefully drafted our IC/EC and in cases where we found them not yet sufficient, we enhanced them. Two experienced researchers carefully read the publications.
Our data source was four famously known databases containing the relevant literature [51]. Due to the extensive snowballing, we made sure that the potential for blind spots, i.e., publications that are only listed by one engine or none at all, is minimized. This also holds for the case of WoS, where different result sizes are obtained. The divergence between both result sets was marginal, so we are confident excessive snowballing alleviates it. The snowballing also alleviated some of our rigorousness by applying IC/EC very early on in our search process. While we reduced our initial tally of 1492 publications to only 89, a reduction of two orders of magnitude, snowballing helped us to ensure recapturing significant contributions into our database again.
@@
Another threat is the use of search terms. We used PICO as Kitchenham et al. [10] suggest to select the appropriate search term. Nevertheless, we know it is impossible to cover all types of FM, AI, and all their acronyms. Therefore, the aim was to conduct an exhausting snowballing procedure to include all relevant publications.
@@
Grey literature was not considered for a variety of reasons. First, we did not encounter large quantities of grey literature during our snowballing. Thus, We can represent current research directions even without considering them. Second, We were concerned about the feasibility of the provided workshop and short papers. As we only did a soft skimming of the content, we can not offer a quality and feasibility assessment of the included contributions. As we aim to provide the reader with only high-quality studies, we decided to require official publications.

### v2

@@ 5.2 RQ 2: Content type of contributions
Lastly, we investigated the data sets that were found. Here, multiple observations took place.
First, we searched for high-quality data sets enabling other researchers to conduct their research efficiently. As pointed out earlier, we found 21 data sets. In Fig. 15, we see the respective scopes of the data sets. Of the 21 data sets, 15 focused on TP, while 3 aimed at synthesis, 1 at SAT, 1 at SMT, and 1 at MC.
However, we were also interested in datasets that were not novel contributions but were found to be reused in other studies. For this, we skimmed all our studies for dataset mentions. As it turns out, of the 189 articles, 124 used external data sets. The 65 remaining studies generate random samples or do not mention their data source. Noteworthy is that when random data is used, the generated data is seldom shared, hindering the reproduction of results.
The datasets found for the 124 contributions are largely heterogeneous. This means that most studies utilized different datasets. However, we could find some data repositories that had multiple usages. The Mizar Library [^6] https://mizar.uwb.edu.pl/library/ (and associated libraries like, e.g., MPTP[46]) saw 23 usages, TPTP Problem Library [^7] https://tptp.org/TPTP/ has 7 mentions, The ProB machine library [^8] https://prob.hhu.de/w/index.php?title=Download got 5 mentions, HoL [^9] https://isabelle.in.tum.de/library/ 5, SMTLib [^10] https://smt-lib.org 5, SATComp [^11] https://satcompetition.github.io 5, SATLib [^12] https://www.cs.ubc.ca/~hoos/SATLIB/benchm.html 4, CoqGym [^13] https://github.com/princeton-vl/CoqGym 4. The full list is available in the repository linked in Section 1.
@@ 6 Discussion
In the following, we will evaluate the results and conclude by answering our research questions. Further, we will outline lessons learned to provide actionable insights for conducting future studies in the area.
@@ 6.1 RQ 1: Demographics of the research area
For RQ 1.1, we see a general growth in the research field. However, from the presented data points, it cannot be inferred alone whether this trend follows the overall trend of AI’s rise in popularity or is independent.
For RQ 1.2, we see a strong focus on conference papers, short papers, and workshops. Together with the singular book chapter, this indicates an emerging field, as there is a tendency to rely on publication methods supporting short review cycles. This is further supported by the fact that the only book chapter was a case study. Furthermore, we see that the AI side of things mainly drives the topic, and especially FM-based conferences publish less content about the application of AI in FM. This could indicate hesitance to adapt or a lack of AI maturity. Contrasting this with RQ 1.1, we assume that the research interest will at least not plummet in the foreseeable future, but might plateau for the time being.
For RQ 1.3, we can see that the bulk of the work lies within the area of methodologies. This could be due to the current paper requirements, which request a novel approach and some evaluation within the paper. Another reason might be AI’s applicative nature, which often allows it to serve as a support system to automate parts of the work. We also want to highlight the possibility that we primarily introduced methodology papers based on the nature of our search query, which focuses on the application of AI in the FM domain. However, due to rigorous snowballing, we are confident we correctly assessed the state of research in this field.
What is more interesting is what we do not see: many case studies and benchmarks. Recalling Fig. 7, we only see two publications providing case studies and two providing benchmarks. This may hint at a further gap in maturity, especially as these contribution types are primarily present in journal and book contributions. One would assume more such publications if a set of robust solutions for everyday problems were readily available so that researchers could compare their effectiveness against each other. It also highlights the need for proper and uniform benchmark sets within the community, which can be shared between different research groups.
@@ 6.2 RQ 2: Content type of contributions
In RQ 2, we were interested in how the AI application is distributed in FM. For this, we posed RQs 2.1, 2.4, 2.2 and 2.3 to investigate which AI techniques find appreciation in the community, which FM areas are already subject to AI application, if certain AI techniques are predominantly used in specific FM fields, and whether there exist publicly available data sets that can be used in further research.
@@
For RQ 2.1, we saw in Fig. 8 that most studies use NNs, followed by RL, and NLP and EAs are far behind. From a paradigmatic point of view, it makes sense to rank these four at the top, as they all attempt to solve different problem settings. NNs are a stand-in for classification and regression tasks, RL is applied for learning behavioral policies, NLP processes texts resembling natural language or code by extension, and EAs excel in optimization problems. However, it is interesting to see NNs being so far ahead of the rest with 83 publications utilizing them. Especially, since they are not the only algorithms suitable for classification and regression tasks, but competing algorithms such as KNN, SVM, or decision trees are, at least as solely applied AI techniques, vastly underrepresented. These algorithms, which we categorized as classical approaches in Section 2.3, are, however, not forgotten, as we can see in publications that contrast multiple different AI algorithms (Fig. 8(b)). This focus on NNs when only one technique is used can have multiple underlying reasons. First, implementing, training, and fine-tuning NNs is time-consuming and might require resources needed to investigate further algorithms. Second, we have seen in Fig. 3 that the research area gained traction in the early 2010s. As this coincides with the advent of deep learning [47], this might suggest the need for NNs to overcome the complexities the area of FM offers, and that more classical ML algorithms are unsuitable for most tasks. Third, the community might be biased towards using NNs from the start instead of applying more straightforward techniques first.
@@
However, the prominence of more classical algorithms in the case of multiple AI techniques shows that the community seems to be aware that other algorithms might be suitable for their problems as well. This is especially amplified by Fig. 9, showing that most frequently, methodologies rely on various techniques to solve problems.
@@
Regarding maturity, Fig. 10 indicates no noteworthy increase in techniques besides NLP. For 2023, we established insufficient data points so that we may see a rise in NLP-based contributions, which would fit a more general advent of NLP-related contributions due to the rise of transformer architectures and LLMs.
@@
For RQ 2.2, we have seen in Fig. 11(a) that TP and SAT are the predominant FM areas in which AI is applied, making up 68.8 % of all publications in our five-year corpus. We explain this by the fact that both areas are also of interest to non-FM communities. A detailed view of the authors’ origin domains would be necessary to verify this assumption. However, this is still beneficial for the FM community as a whole. For instance, Fig. 12(a) shows a strong focus on FM-relevant daily use cases. Figure 12(b) reveals much current work regarding first-order logic on the especially relevant FM topics of premise selection. While this trend is good, complex FM models rely on complicated data types and require sophisticated reasoning.
@@
Regarding SAT, we can take away from Fig. 11(c) that much focus lies directly on SAT solving. At the same time, other authors also attempt to predict satisfiability, i.e., whether a formula is SAT or UNSAT. The latter seems to contradict the rigorousness needed in FM, as the probabilistic nature of AI approaches lacks the certainty of the predicted satisfiability. This can be a product of cross-fertilization, as SAT is also relevant for non-FM communities. Here, a reasonable estimate about a formula’s SAT or UNSAT may be more appreciated than in the FM community.
@@
Figure 11(d) shows that the interest in synthesis is low compared to other topics and primarily focuses on model generation. Given the recent rise of generative AI approaches, we see room for promising research in this area.
@@
A similar low interest seems to be in model checking. Figure 13(b) shows a relatively low number of contributions in that area. The direct comparison with the TP field is interesting, as MC also contributes very little to tool and tool contributions, hinting at less application-focused research. Figure 13(a) shows a relatively consistent interest in recent years, while Fig. 13(b) emphasizes how little is done for current practitioners in terms of tool (improvements), benchmarks, training data, and case studies.
@@
For RQ 2.3, we saw in Fig. 14(a) that NNs are the undeniable favorite in almost all FM areas investigated. We already discussed why NNs might be more popular than other algorithms suitable for the same problem tasks above for RQ 2.1. However, this again suggests that the FM community mainly focuses on solving classification or regression problems rather than execution policies (RL) or direct processing of texts or code (NLP). The slight exception we see is in model checking, which focuses mainly on EAs, which interpret model checking as a sort of optimization problem, i.e., finding a path to a counter-example. From building the larger corpus of 457 studies, we know from seeing various titles that we missed many EA papers due to our 5-year restriction. Presumably, EA was more prominently used before deep learning took off.
@@
Finally, for RQ 2.4, we found 19 datasets (or training environment) contributions in our corpus of 189 studies. Considering the two separate, data set-only publications that did not pass IC 1 (see Section 4.5), this leaves us with 21 total data set contributions.
With the analyzed high heterogeneity of the used datasets, as already mentioned in Section 5.1.3, there may be a general underlying issue. There seems to be no defined benchmark gold standard against which to run new insights. Instead, results are run on the data sets at hand. Furthermore, while datasets are publicly available for TP and SAT solving, the number of usable, established, and available benchmark datasets for other research areas that fit the named criteria is minuscule.
@@
In conclusion, we strongly need unified defaults to allow reproducibility and introduce familiar benchmarks into the field. Possible issues hindering such a development might be found in the various existing formalisms, which are not necessarily used or even supported by individual research groups. Such boundaries, of course, weaken the impact a data set publication might have onto researchers working with differing formalisms.
6.3 Observed objectives of AI applications in FM
Generally, we see that the stark majority of publications use AI as an assistance tool to enable or enhance a respective FM tool’s performance. That is, AI is used to find models, clauses, or lemmas. However, the found artifacts are then still checked by a formal tool.
With the notable exception of SAT solving, where 5 contributions directly predict the satisfiability of a formula with AI, we did not observe any focus on finding AI-generated guarantees for formal models to replace an FM tool. This seems only natural, as FM are about believable and reproducible guarantees. Therefore, giving most AI approaches a more prominent role comes with the price of admission: introducing determinism and uncertainty. For strictly formal-safety-related projects, this may be too much of a burden to pass.
Another observation we did not make was the presence of auto-active verification. Analyzing existing program code did not draw much attention, and the only remotely connected contribution was outside our search area.
@@ 6.4 Suggested research directions
As pointed out in our discussion to RQ 2.4, we are concerned with the lack of established data sets and benchmarks in the FM sub-communities. While many contributions have already tackled this problem for TP, the other areas seem to fall short. This creates room for reproducibility issues between research groups and hinders a more precise comparison between results and established benchmarks that could be provided.
Developing strong baselines might increase comparability between different methods and reduce the entry hurdle for new researchers due to publicly available data sets. The main challenge is to create data sets that can be used for different formalisms. This would allow smaller communities to still benefit from the available data, which might be presented in a more popular formalism while furthering research in their area.
An orientation on how such a unified benchmark environment could look would be OpenAIGym [49], a Toolkit for RL to foster research by enhancing reproducibility.
@@
In the age of big data, we were expecting more studies that apply data mining techniques, be it for feature engineering or performance analytical approaches. This was not the case. While it might very well be that the kind of problems solvable by data mining are irrelevant to the FM communities, it is also possible that potential benefits are not well-known.
We propose to research whether and how data mining approaches can benefit FM. First, this would highlight potential benefits to the community. Second, due to the investigative nature of data mining, this might strengthen our insights into our tools and best practices. Consequently, this might further the quality of other AI applications in FM due to new knowledge.
@@
While LLMs, at this point, are established tools that have found their way into daily applications, we have found little use in our corpus of studies. This can be due to their relatively high employment costs. Another reason could be that respective studies are still published and invisible to our systematic mapping study. Nevertheless, we see strong potential for generative AI approaches, especially in synthesis and auto-formalization. However, simple test cases or documentation generation for existing models seems a promising starting point for building trust.
We envision substantial benefits from applying generative AI during FM development and highlight the need for more research. One can easily envision tool chains of automatic generation, checking, and documentation of formal models, logical formulas, or proofs with reproducible and provable results. However, we also see the need for more exploration regarding how generative AI can be helpful besides text production. For instance, how could generative AI be leveraged in proof mining? What stages in the formal development process can benefit from generated inputs, and what types could and should be generated? Generative AI for FM may be thought of further than pure text and code generation and more toward reasoning support.
Increase the potential of AI in model checking
Another underrepresented topic is model checking. Anecdotally, we know model checking had much EA applied in the past, while it was a seemingly unpopular application area for AI in our five-year corpus. However, we see significant potential in revisiting the research with a modern lens. Especially in the context of bounded model checking (BMC) or statistical model checking (SMC) and faulty state prediction. BMC cuts off parts of a model’s state space for performance and computability. SMC checks states until a certain confidence in the absence of faulty states is reached. Both approaches do not fully traverse the state space and can only indicate the absence of faults.
AI-enforced model checking could complement BMC, SMC, and model checking by predicting potential faulty states or needed search depths, learning search policies via RL to reach faulty states more targeted, or predicting equivalence classes of states for search space reduction. For example, after estimating a defective state or a search depth, the model checker could be applied in a very targeted manner to confirm the predicted state.
@@ 7 Threats to validity
A study of this type and extent can only be conducted with the possibility of encountering threats that threaten validity. Zhou et al. [30] serves as a baseline for this discussion, as their contribution meticulously lists the possible threats to validity and how they may be addressed.
@@
A potential threat to the conclusion could be our selection of articles and our data source. To ensure the quality of our data, we carefully drafted our IC/EC, and in cases where we found them not yet sufficient, we enhanced them. Two experienced researchers, i.e., the first two authors, carefully read the publications.
Our data source was four famous known databases containing the relevant literature [50]. Due to the extensive snowballing, we made sure that the potential for blind spots, i.e., publications that are only listed by one engine or none at all, is minimized. This also holds for the case of WoS, where different result sizes are obtained. The divergence between the two results was marginal, so we are confident that excessive snowballing alleviates it. The snowballing also alleviated some of our rigor by applying IC/EC early in our search process. While we reduced our initial tally of 1492 publications to only 89, a reduction of two orders of magnitude, snowballing helped us recapture significant contributions into our database again.
@@
Another threat is the selection and use of search terms. We used PICO as Kitchenham et al. [10] suggest to select the appropriate search term. Nevertheless, we know it is impossible to cover all types of FM, AI, and all their acronyms in one query due to the technical limitations of the search engines and the practical feasibility of thoroughly skimming the amount of papers found this way. Therefore, the aim was to keep the initial set small and conduct an exhaustive snowballing procedure to include all relevant publications.
@@
While we followed Kitchenham et al. [10] guidelines for a successful mapping study to minimize the threat to the conclusion of this study, one remaining potential threat is whether or not to include grey literature, as we could have missed crucial contributions that change the outcome of this study. Grey literature was not considered for two main reasons. First, we did not encounter large quantities of grey literature during our snowballing. Thus, We are positive that we can represent current research directions without considering them. Second, we only did a soft skim of the content, and therefore, we can not offer a quality and feasibility assessment of the included contributions. As we aim to provide the reader with only high-quality studies, we decided to require official publications.